# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 6.2. Comparison against a hull
On the main branch in the iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ all extenders used are fully backgrounded, i.e., for every $E_{\alpha}^{\mathcal{T}}$, $\alpha\in[0,\lambda]_{T}$, and every $l<\omega$,
$\operatorname{crit}(E_{\alpha}^{\mathcal{T}})\neq i_{0\alpha}^{\mathcal{T}}(\kappa_{l}).$
As the construction of $\mathcal{M}$ did not move up to this point of the comparison, this means that disagreements were only caused by fully backgrounded extenders. All extenders used in the iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ have critical point above $\kappa_{i}$ as ¯$\bar{\mathcal{M}}|\kappa_{i}=\mathcal{M}|\kappa_{i}$ and none of the backgrounded extenders overlaps $\kappa_{i}$. Write $\kappa_{k}^{*}$ for the image of $\bar{\kappa}_{k}$ in $\mathcal{M}^{*}$ for all $k<\omega$ and note that $\kappa_{k}^{*}=\kappa_{k}$ for all $k\leq i$. Recall that $\operatorname{crit}(E)=\kappa_{j}<\kappa_{i}$, so the extender $F$ on $\mathcal{M}^{*}$ has critical point $\kappa_{j}^{*}=\kappa_{j}$.
We can at the same time consider the iterate $\mathcal{W}^{*}$ of $\bar{\mathcal{W}}$ via the background extenders and maintain the property that $\mathcal{M}^{*}$ is the result of an $\mathcal{M}$-construction inside $\mathcal{W}^{*}$. In particular, $F$ is generically countably complete and hence by Lemma 5.17 certified in $\mathcal{W}^{*}$ as it has critical point $\kappa_{j}=\kappa_{j}^{*}<\kappa_{i}$ and index above $\kappa_{i}$.
By [12, Theorem 3.3], inside $\mathcal{W}$, each $\mathcal{P}^{i}$ has a unique iteration strategy $\Sigma_{i}$. So, since $\bar{\mathcal{W}}$ is a hull inside $\mathcal{W}$, $\bar{\Sigma}_{i}$ and $\Sigma_{i}$ agree in $\mathcal{W}$ for all $i<\omega$. The $\bar{\mathcal{W}}$-to-$\mathcal{W}^{*}$ iteration can also be performed inside $\mathcal{W}$ and by strategy coherence in $\mathcal{W}$, the tail strategy $\Sigma_{i}^{*}$ of $\bar{\Sigma}_{i}$ along the $\bar{\mathcal{W}}$-to-$\mathcal{W}^{*}$ iteration is equal to $\Sigma_{i}$ for all $i<\omega$. Strategy coherence holds for $\Sigma_{i}$ by the argument [^19] This argument works for Mitchell-Steel indexing without the additional fine structural assumptions made in [27]. We thank Benjamin Siskind for pointing this out. for [27, Corollary 7.6.9] or as a consequence of positionality, which holds for $\Sigma_{i}$ as shown in [12]. Therefore, not only $\mathcal{M}^{*}$ and $\mathcal{M}$ agree below the index of $F$ but also the strategies of their backgrounds $\mathcal{W}^{*}$ and $\mathcal{W}$ on $\mathcal{P}^{j}$. So, in particular, they have the same tail strategy $(\Sigma_{j})_{\mathcal{M}_{\infty}^{j,-}}$ of $\Sigma_{j}$. As $E$ and $F$ are both certified via this strategy, this implies $E=F$.
@@
The iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ uses in addition to fully backgrounded extenders a single extender $G$ that is not fully backgrounded.
We again aim to show that $F$ is certified via the same tail strategy of $\Sigma_{j}$ that certifies $E$. For notational simplicity we suppose that the iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ only uses $G$, i.e., $\mathcal{M}^{*}=\operatorname{Ult}(\bar{\mathcal{M}},G)$. Taking additional iterations via fully backgrounded extenders into account is a straightforward generalization of the argument below so we leave that to the reader. Write $(\kappa_{i}^{*},\delta_{i}^{*}\mid i<\omega)$ for the image of $(\kappa_{i},\delta_{i}\mid i<\omega)$ under $\pi_{G}\colon\bar{\mathcal{M}}\rightarrow\operatorname{Ult}(\bar{\mathcal{M}},G)$.
@@
This is a straightforward generalization of the previous case using the fact that we can still find realization maps for iterates of countable substructures as in the iterability proof, cf. Lemma 6.3.
@@
Let $(E_{l}\mid l<\gamma)$ be the sequence of extenders used that have critical point the image of some $\bar{\kappa}_{l}$ for $l<\omega$. Recall that the extenders need to be of increasing length and note that if there is some $l<k<\gamma$ such that $\operatorname{crit}(E_{l})>\operatorname{crit}(E_{k})$ and $E_{k}$ is the first extender where this happens after stage $l$, then $E_{k}$ gets applied to $\bar{\mathcal{M}}$. As we are only interested in the extenders on the main branch leading from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$, we might as well assume that the sequence $(\operatorname{crit}(E_{l})\mid l<\gamma)$ is nondecreasing.
@@
Note that even in the case that $\operatorname{crit}(F)=\kappa_{k}^{*}$ we can define the direct limit system that gives rise to the direct limit $(\mathcal{M}_{\infty}^{k})^{*}$ to characterize $F$ in $\mathcal{M}^{*}$ as in Section 5 via the internalized limit of $\mathcal{N}_{\delta^{*}}$’s for $\kappa_{k}^{*}$-weak cutpoints $\delta^{*}$ in $\mathcal{M}^{*}|\kappa_{k}^{*}$ using elementarity between $\bar{\mathcal{M}}$ and $\mathcal{M}^{*}$. By agreement in the comparison process, $\mathcal{M}^{*}|\kappa_{k}^{*}=\mathcal{M}|\kappa_{k}$ and $(\mathcal{M}_{\infty}^{k})^{*}=\mathcal{M}_{\infty}^{k}$.
@@
Extender only on the $\mathcal{M}$ side: We finally argue that the reason $E$ is used on the $\mathcal{M}$-side of the comparison cannot be that there is no extender $F$ with the same index as $E$ on $\mathcal{M}^{*}$. So say $E=E_{\alpha}^{\mathcal{M}}$ and suppose toward a contradiction that $E_{\alpha}^{\mathcal{M}^{*}}=\emptyset$. Recall that $\operatorname{crit}(E_{\alpha}^{\mathcal{M}})=\kappa_{j}\leq\kappa_{i}$.
@@
The iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ does not use any extenders with critical point ${\leq}\kappa_{j}$ on the main branch.
@@
Suppose toward a contradiction that there is some extender $F$ with critical point $\kappa_{l}\leq\kappa_{j}$ that gets used on the main branch in the iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$.
@@
In case $\kappa_{j}<\kappa_{i}$, recall that we suppose that $\kappa_{l}$ is fully strong in $\mathcal{M}$. So there is some extender $E_{\beta}^{\mathcal{M}}$ on the sequence of $\mathcal{M}$ with critical point $\kappa_{l}$ and index $\beta$ above the index of $F$ on the $\mathcal{M}^{*}$-sequence. If $\kappa_{l}=\kappa_{j}=\kappa_{i}$, $F$ is indexed on the $\mathcal{M}^{*}$-sequence below $\delta_{i+1}$ as $\mathcal{M}^{*}\cap\operatorname{Ord}<\delta_{i+1}$. Since $\kappa_{i}$ is ${<}\delta_{i+1}$-strong in $\mathcal{M}$, there is an extender $E_{\beta}^{\mathcal{M}}$ on the $\mathcal{M}$-sequence with critical point $\kappa_{i}$ and $\operatorname{index}^{\mathcal{M}}(E_{\beta}^{\mathcal{M}})>\operatorname{index}^{\mathcal{M}^{*}}(F)$. In both cases, let $\beta$ be minimal with this property and write $\alpha$ for the $\mathcal{M}^{*}$-index of $F$. Then $\alpha$ is a cutpoint in $\operatorname{Ult}(\mathcal{M},E_{\beta}^{\mathcal{M}})$ and the $(\kappa_{l},\alpha)$-extender $E$ derived from the ultrapower embedding $\pi^{\mathcal{M}}_{E_{\beta}^{\mathcal{M}}}$ is on the $\mathcal{M}$-sequence by the initial segment condition. So $E_{\alpha}^{\mathcal{M}}=E$ and $E$ is on the sequence of $\mathcal{M}_{\nu}$ for some $\nu\geq\zeta$. Therefore, even if $F$ is used in the comparison against $\mathcal{M}_{\zeta}$ it will not be used on the main branch in the comparison of $\bar{\mathcal{M}}$ against the construction of $\mathcal{M}$ as this part of the coiteration gets collapsed when moving on to the comparison with $\mathcal{M}_{\nu}$. So we might and will assume that the iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ involves comparisons against $\mathcal{M}_{\nu}$ or larger models from the construction of $\mathcal{M}$.
@@
Now consider the case $\kappa_{j}=\kappa_{i}$. In this case we use a more complicated argument to derive a contradiction. Let, as before, $F$ be the extender causing the disagreement on the $\mathcal{M}$-side or any long enough extender with critical point $\kappa_{i}$ on the model $\mathcal{M}_{\zeta}$ in the construction of $\mathcal{M}$ that is currently compared against $\mathcal{M}^{*}$. Consider the extender $E$ with critical point $\kappa_{i}$ given by
$(a,A)\in E\text{ iff }(i(a),A)\in F$
for $a\in[\operatorname{Ord}]^{<\omega}$ and $A\subseteq[\kappa_{i}]^{|a|}$. By the last subclaim, ¯$\bar{\mathcal{M}}\cap\mathcal{P}(\kappa_{i})=\mathcal{M}^{*}\cap\mathcal{P}(\kappa_{i})=\mathcal{M}_{\zeta}\cap\mathcal{P}(\kappa_{i})$ and this defines an extender $E$ over $\bar{\mathcal{M}}$. Let $\bar{\xi}$ be the least $i$-weak cutpoint above $\kappa_{i}$ in $\operatorname{Ult}(\bar{\mathcal{M}},E)$. Recall that an ordinal $\alpha$ is an $i$-weak cutpoint above $\kappa_{i}$ if all extenders overlapping $\alpha$ have critical point $\kappa_{j}$ for some $j<i$. Then $\bar{\eta}=(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},E)}$ is the ordinal where $E$ would be indexed if it is on the sequence of $\bar{\mathcal{M}}$.
We argue that $E$ coheres ¯$\bar{\mathcal{M}}|\bar{\eta}$ and is generically countably complete in $\mathcal{W}$. Consider the commuting diagram in Figure 8, where $\sigma$ is given by
$[a,f]_{E}\mapsto[i(a),i(f)]_{F}.$
$\mathcal{M}_{\zeta}|i(\bar{\xi})$$j_{F}$$j_{F}(\mathcal{M}_{\zeta}|i(\bar{\xi}))$¯$\bar{\mathcal{M}}|\bar{\xi}$$j_{E}$$j_{E}(\bar{\mathcal{M}}|\bar{\xi})$$i$$\sigma$
@@
For $a\in[\bar{\eta}]^{<\omega}$, we have $a=[a,\operatorname{id}]_{E}^{\bar{\mathcal{M}}}$, so in particular for all ordinals $\chi\leq\bar{\xi}$,
$\sigma(\chi)=\sigma([\chi,\operatorname{id}]_{E}^{\bar{\mathcal{M}}})=[i(\chi),\operatorname{id}]_{F}^{\mathcal{M}_{\zeta}}=i(\chi).$
@@
¯$\bar{\mathcal{M}}|\bar{\xi}=\operatorname{Ult}(\bar{\mathcal{M}},E)|\bar{\xi}.$
@@
Let $X=\sigma\,"(\operatorname{Ult}(\bar{\mathcal{M}},E)|\bar{\xi})$ and $Y=i\,"(\bar{\mathcal{M}}|\bar{\xi})$. We will argue that $X=Y$ using that they are both elementary substructures of the same model and have the same ordinals.
First, note that $X$ and $Y$ are both (uncollapsed) elementary substructures of $\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$ as $\mathcal{M}_{\zeta}|i(\bar{\xi})=\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$ using that $F$ coheres $\mathcal{M}_{\zeta}|\eta$.
$\operatorname{Ult}(\mathcal{M}_{\zeta},F)$ is a fine structural premouse, so every set in $\operatorname{Ult}(\mathcal{M}_{\zeta},F)$ is ordinal definable (with the extender sequence as a predicate). Let $x\in X$ be arbitrary and let $\varphi_{x}(v,v_{0},\dots,v_{n})$ be a formula defining $x$ in $\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$. Then
$\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})\vDash\exists\alpha_{0},\dots,\alpha_{n}\,\varphi_{x}(x,\alpha_{0},\dots,\alpha_{n})\wedge\forall z(\varphi_{x}(z,\alpha_{0},\dots,\alpha_{n})\rightarrow x=z).$
@@
As $X\cap\operatorname{Ord}=Y\cap\operatorname{Ord}$, in particular $\xi_{0},\dots,\xi_{n}\in Y$ and we have by elementarity between $X$ and $\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$ as well as between $Y$ and $\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$,
@@
As $Y\preceq\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$ and $x$ is the unique element in $\operatorname{Ult}(\mathcal{M}_{\zeta},F)|i(\bar{\xi})$ satisfying $\varphi_{x}(x,\xi_{0},\dots,\xi_{n})$, this in fact implies $x\in Y$. So we have shown that $X\subseteq Y$. The same argument shows $Y\subseteq X$ and we have $X=Y$, as desired. ∎
The subclaim implies ¯$\bar{\mathcal{M}}|\bar{\xi}\unlhd j_{E}(\bar{\mathcal{M}}|(\kappa_{i}^{+})^{\bar{\mathcal{M}}})$. Recall that for coherence we aim to show that in fact
¯$\bar{\mathcal{M}}|(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},E)}=j_{E}(\bar{\mathcal{M}}|(\kappa_{i}^{+})^{\bar{\mathcal{M}}})|(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},E)}.$
We will use the operators for this argument. In the following subclaim we show that $E$ is given by operators and hence is the extender derived from the hull embedding $k\colon S_{i}(\bar{\mathcal{M}}|\kappa_{i})\rightarrow S_{i}(\bar{\mathcal{M}}|\bar{\xi})$.

### v3

@@ 6.2. Comparison against a hull
On the main branch in the iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ all extenders used are fully backgrounded, i.e., for every $E_{\alpha}^{\mathcal{T}}$, $\alpha\in[0,\lambda]_{T}$, and every $l<\omega$,
$\operatorname{crit}(E_{\alpha}^{\mathcal{T}})\neq i_{0\alpha}^{\mathcal{T}}(\kappa_{l}^{\bar{\mathcal{M}}_{\xi}}).$
As the construction of $\mathcal{M}$ did not move up to this point of the comparison, this means that disagreements were only caused by fully backgrounded extenders. All extenders used in the iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ have critical point above $\kappa_{i}$ as ¯$\bar{\mathcal{M}}|\kappa_{i}=\mathcal{M}|\kappa_{i}$ and none of the backgrounded extenders overlaps $\kappa_{i}$. Write $\kappa_{k}^{*}$ for the image of $\bar{\kappa}_{k}$ in $\mathcal{M}^{*}$ for all $k<\omega$ and note that $\kappa_{k}^{*}=\kappa_{k}$ for all $k\leq i$. Recall that $\operatorname{crit}(E)=\kappa_{j}<\kappa_{i}$, so the extender $F$ on $\mathcal{M}^{*}$ has critical point $\kappa_{j}^{*}=\kappa_{j}$.
Since we are comparing the constructions of $\bar{\mathcal{M}}$ and $\mathcal{M}$, we can at the same time consider the iterate $\mathcal{W}^{*}$ of $\bar{\mathcal{W}}$ via the background extenders and maintain the property that $\mathcal{M}^{*}$ is the result of an $\mathcal{M}$-construction inside $\mathcal{W}^{*}$. In particular, $F$ is generically countably complete and hence by Lemma 5.17 certified in $\mathcal{W}^{*}$ as it has critical point $\kappa_{j}=\kappa_{j}^{*}<\kappa_{i}$ and index above $\kappa_{i}$.
By [13, Theorem 3.3], inside $\mathcal{W}$, each $\mathcal{P}^{i}$ has a unique iteration strategy $\Sigma_{i}$. So, since $\bar{\mathcal{W}}$ is a hull inside $\mathcal{W}$, $\bar{\Sigma}_{i}$ and $\Sigma_{i}$ agree in $\mathcal{W}$ for all $i<\omega$. The $\bar{\mathcal{W}}$-to-$\mathcal{W}^{*}$ iteration can also be performed inside $\mathcal{W}$ and by strategy coherence in $\mathcal{W}$, the tail strategy $\Sigma_{i}^{*}$ of $\bar{\Sigma}_{i}$ along the $\bar{\mathcal{W}}$-to-$\mathcal{W}^{*}$ iteration is equal to $\Sigma_{i}$ for all $i<\omega$. Strategy coherence holds for $\Sigma_{i}$ by the argument [^33] This argument works for Mitchell-Steel indexing without the additional fine structural assumptions made in [29]. We thank Benjamin Siskind for pointing this out. for [29, Corollary 7.6.9] or as a consequence of positionality, which holds for $\Sigma_{i}$ as shown in [13]. Therefore, not only $\mathcal{M}^{*}$ and $\mathcal{M}$ agree below the index of $F$ but also the strategies of their backgrounds $\mathcal{W}^{*}$ and $\mathcal{W}$ on $\mathcal{P}^{j}$. So, in particular, they have the same tail strategy $(\Sigma_{j})_{\mathcal{M}_{\infty}^{j,-}}$ of $\Sigma_{j}$. As $E$ and $F$ are both certified via this strategy, this implies $E=F$.
@@
The iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ uses in addition to fully backgrounded extenders a single extender $G$ that is not fully backgrounded. [^34] This case might actually occur for extenders $G$ with critical point the image of some $\bar{\kappa}_{l}$ for $l>i$.
We again aim to show that $F$ is certified via the same tail strategy of $\Sigma_{j}$ that certifies $E$ and hence $E$ does not cause a disagreement in the comparison. For notational simplicity we suppose that the iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ only uses $G$, i.e., $\mathcal{M}^{*}=\operatorname{Ult}(\bar{\mathcal{M}}_{\xi},G)$. Taking additional iterations via fully backgrounded extenders into account is a straightforward generalization of the argument below so we leave that to the reader. Write $(\kappa_{i}^{*},\delta_{i}^{*}\mid i<\omega)$ for the image of $(\bar{\kappa}_{i},\bar{\delta}_{i}\mid i<\omega)$ under $\pi_{G}\colon\bar{\mathcal{M}}_{\xi}\rightarrow\operatorname{Ult}(\bar{\mathcal{M}}_{\xi},G)$, if these cardinals are defined in ¯$\bar{\mathcal{M}}_{\xi}$.
@@
This is a straightforward generalization of the previous case using the fact that we can still find realization maps for iterates of countable substructures as in the iterability proof, cf. Lemma 6.3. Therefore, we also get in Case 3 that $E$ does not cause a disagreement.
@@
Let $(E_{l}\mid l<\gamma)$ be the sequence of extenders used that have critical point the image of some $\bar{\kappa}_{l}$ for $l<\omega$. Recall that the extenders need to be of increasing length and note that if there is some $l<k<\gamma$ such that $\operatorname{crit}(E_{l})>\operatorname{crit}(E_{k})$ and $E_{k}$ is the first extender where this happens after stage $l$, then $E_{k}$ gets applied to ¯$\bar{\mathcal{M}}_{\xi}$. As we are only interested in the extenders on the main branch leading from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$, we might as well assume that the sequence $(\operatorname{crit}(E_{l})\mid l<\gamma)$ is nondecreasing.
@@
Note that even in the case that $\operatorname{crit}(F)=\kappa_{k}^{*}$ we can define the direct limit system that gives rise to the direct limit $(\mathcal{M}_{\infty}^{k})^{*}$ to characterize $F$ in $\mathcal{M}^{*}$ as in Section 5 via the internalized limit of $\mathcal{N}_{\delta^{*}}$’s for $\kappa_{k}^{*}$-weak cutpoints $\delta^{*}$ in $\mathcal{M}^{*}|\kappa_{k}^{*}$ using elementarity between ¯$\bar{\mathcal{M}}_{\xi}$ and $\mathcal{M}^{*}$. By agreement in the comparison process, $\mathcal{M}^{*}|\kappa_{k}^{*}=\mathcal{M}|\kappa_{k}$ and $(\mathcal{M}_{\infty}^{k})^{*}=\mathcal{M}_{\infty}^{k}$.
@@
This finishes the argument that $E$ does not cause a disagreement in the case that there is an extender $F$ with the same index in the $\bar{\mathcal{M}}$-side of the comparison.
Extender only on the $\mathcal{M}$ side: We finally argue that the reason $E$ is used on the $\mathcal{M}$-side of the comparison cannot be that there is no extender $F$ with the same index as $E$ on $\mathcal{M}^{*}$. So say $E=E_{\alpha}^{\mathcal{M}}$ and suppose toward a contradiction that $E_{\alpha}^{\mathcal{M}^{*}}=\emptyset$. Recall that $\operatorname{crit}(E_{\alpha}^{\mathcal{M}})=\kappa_{j}\leq\kappa_{i}$. We start by showing that the iteration on the $\bar{\mathcal{M}}$-side does not use any extenders with critical point $\leq\kappa_{j}$.
@@
The iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ does not use any extenders with critical point ${\leq}\kappa_{j}$.
@@
Suppose toward a contradiction that there is some extender $F$ with critical point the image of $\kappa_{l}=\bar{\kappa}_{l}\leq\bar{\kappa}_{j}=\kappa_{j}$ that gets used in the iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$.
@@
In case $\kappa_{j}<\kappa_{i}$, recall that we suppose that $\kappa_{l}$ is fully strong in $\mathcal{M}$. So there is some extender $E_{\beta}^{\mathcal{M}}$ on the sequence of $\mathcal{M}$ with critical point $\kappa_{l}$ and index $\beta$ above the index of $F$ on the $\mathcal{M}^{*}$-sequence. If $\kappa_{l}=\kappa_{j}=\kappa_{i}$, $F$ is indexed on the $\mathcal{M}^{*}$-sequence below $\delta_{i+1}$ as $\mathcal{M}^{*}\cap\operatorname{Ord}<\delta_{i+1}$. Since $\kappa_{i}$ is ${<}\delta_{i+1}$-strong in $\mathcal{M}$, there is an extender $E_{\beta}^{\mathcal{M}}$ on the $\mathcal{M}$-sequence with critical point $\kappa_{i}$ and $\operatorname{index}^{\mathcal{M}}(E_{\beta}^{\mathcal{M}})>\operatorname{index}^{\mathcal{M}^{*}}(F)$. In both cases, let $\beta$ be minimal with this property and write $\alpha$ for the $\mathcal{M}^{*}$-index of $F$. Then $\alpha$ is a cutpoint in $\operatorname{Ult}(\mathcal{M},E_{\beta}^{\mathcal{M}})$ and the $(\kappa_{l},\alpha)$-extender $E$ derived from the ultrapower embedding $\pi^{\mathcal{M}}_{E_{\beta}^{\mathcal{M}}}$ is on the $\mathcal{M}$-sequence by the initial segment condition. So $E_{\alpha}^{\mathcal{M}}=E$ and $E$ is on the sequence of $\mathcal{M}_{\gamma}$ for some $\gamma\geq\zeta$. Therefore, even if $F$ is used in the comparison against $\mathcal{M}_{\zeta}$ it will not be used in the iteration of ¯$\bar{\mathcal{M}}_{\xi}$ into the construction of $\mathcal{M}$ as this part of the comparison gets collapsed when moving on to the comparison with $\mathcal{M}_{\gamma}$. So we might and will assume that the iteration from ¯$\bar{\mathcal{M}}_{\xi}$ to $\mathcal{M}^{*}$ involves comparisons against $\mathcal{M}_{\gamma}$ or larger models from the construction of $\mathcal{M}$.
@@
We will now argue that there is an extender $F$ with the same index as $E$ on $\mathcal{M}^{*}$. For notational simplicity, we write ¯$\bar{\mathcal{M}}_{\xi}=\bar{\mathcal{M}}$ in what follows.
@@
Now consider the case $\kappa_{j}=\kappa_{i}$. In this case we use a more complicated argument to derive a contradiction. Let, as before, $E$ be the extender causing the disagreement on the $\mathcal{M}$-side or any long enough extender with critical point $\kappa_{i}$ on the model $\mathcal{M}_{\zeta}$ in the construction of $\mathcal{M}$ that is currently compared against $\mathcal{M}^{*}$. Consider the extender $F$ with critical point $\kappa_{i}$ given by
$(a,A)\in F\text{ iff }(i(a),A)\in E$
for $a\in[\operatorname{Ord}]^{<\omega}$ and $A\subseteq[\kappa_{i}]^{|a|}$. By the last subclaim, ¯$\bar{\mathcal{M}}\cap\mathcal{P}(\kappa_{i})=\mathcal{M}^{*}\cap\mathcal{P}(\kappa_{i})=\mathcal{M}_{\zeta}\cap\mathcal{P}(\kappa_{i})$ and this defines an extender $F$ over $\bar{\mathcal{M}}$. Let $\bar{\xi}$ be the least $i$-weak cutpoint above $\kappa_{i}$ in $\operatorname{Ult}(\bar{\mathcal{M}},F)$. Recall that an ordinal $\alpha$ is an $i$-weak cutpoint above $\kappa_{i}$ if all extenders overlapping $\alpha$ have critical point $\kappa_{j}$ for some $j<i$. Then $\bar{\eta}=(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},F)}$ is the ordinal where $F$ would be indexed if it is on the sequence of $\bar{\mathcal{M}}$.
We argue that $F$ coheres ¯$\bar{\mathcal{M}}|\bar{\eta}$ and is generically countably complete in $\mathcal{W}$. Consider the commuting diagram in Figure 8, where $\sigma$ is given by
$[a,f]_{F}\mapsto[i(a),i(f)]_{E}.$
$\mathcal{M}_{\zeta}|i(\bar{\xi})$$j_{E}$$j_{E}(\mathcal{M}_{\zeta}|i(\bar{\xi}))$¯$\bar{\mathcal{M}}|\bar{\xi}$$j_{F}$$j_{F}(\bar{\mathcal{M}}|\bar{\xi})$$i$$\sigma$
@@
For $a\in[\bar{\eta}]^{<\omega}$, we have $a=[a,\operatorname{id}]_{F}^{\bar{\mathcal{M}}}$, so in particular for all ordinals $\chi\leq\bar{\xi}$,
$\sigma(\chi)=\sigma([\chi,\operatorname{id}]_{F}^{\bar{\mathcal{M}}})=[i(\chi),\operatorname{id}]_{E}^{\mathcal{M}_{\zeta}}=i(\chi).$
@@
¯$\bar{\mathcal{M}}|\bar{\xi}=\operatorname{Ult}(\bar{\mathcal{M}},F)|\bar{\xi}.$
@@
Let $X=\sigma\,"(\operatorname{Ult}(\bar{\mathcal{M}},F)|\bar{\xi})$ and $Y=i\,"(\bar{\mathcal{M}}|\bar{\xi})$. We will argue that $X=Y$ using that they are both elementary substructures of the same model and have the same ordinals.
First, note that $X$ and $Y$ are both (uncollapsed) elementary substructures of $\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$ as $\mathcal{M}_{\zeta}|i(\bar{\xi})=\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$ using that $E$ coheres $\mathcal{M}_{\zeta}|\eta$.
$\operatorname{Ult}(\mathcal{M}_{\zeta},E)$ is a fine structural premouse, so every set in $\operatorname{Ult}(\mathcal{M}_{\zeta},E)$ is ordinal definable (with the extender sequence as a predicate). Let $x\in X$ be arbitrary and let $\varphi_{x}(v,v_{0},\dots,v_{n})$ be a formula defining $x$ in $\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$. Then
$\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})\vDash\exists\alpha_{0},\dots,\alpha_{n}\,\varphi_{x}(x,\alpha_{0},\dots,\alpha_{n})\wedge\forall z(\varphi_{x}(z,\alpha_{0},\dots,\alpha_{n})\rightarrow x=z).$
@@
As $X\cap\operatorname{Ord}=Y\cap\operatorname{Ord}$, in particular $\xi_{0},\dots,\xi_{n}\in Y$ and we have by elementarity between $X$ and $\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$ as well as between $Y$ and $\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$,
@@
As $Y\preceq\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$ and $x$ is the unique element in $\operatorname{Ult}(\mathcal{M}_{\zeta},E)|i(\bar{\xi})$ satisfying $\varphi_{x}(x,\xi_{0},\dots,\xi_{n})$, this in fact implies $x\in Y$. So we have shown that $X\subseteq Y$. The same argument shows $Y\subseteq X$ and we have $X=Y$, as desired. ∎
The subclaim implies ¯$\bar{\mathcal{M}}|\bar{\xi}\unlhd j_{F}(\bar{\mathcal{M}}|(\kappa_{i}^{+})^{\bar{\mathcal{M}}})$. Recall that for coherence we aim to show that in fact
¯$\bar{\mathcal{M}}|(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},F)}=j_{F}(\bar{\mathcal{M}}|(\kappa_{i}^{+})^{\bar{\mathcal{M}}})|(\bar{\xi}^{+})^{\operatorname{Ult}(\bar{\mathcal{M}},F)}.$
We will use the operators for this argument. In the following subclaim we show that $F$ is given by operators and hence is the extender derived from the hull embedding $k\colon S_{i}(\bar{\mathcal{M}}|\kappa_{i})\rightarrow S_{i}(\bar{\mathcal{M}}|\bar{\xi})$.